# Do More Suspicious Transaction Reports Lead to More Convictions for Money Laundering?

### v1

@@ 5.1 Pooled Regression Results
Table 4: Pooled regression results with $t$ (i.e., “time”) as a regressor. The table contains results for multiple model specifications, numbered P1T to P4T, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level; and *** at the $0.001$ $p$-level.
Model ID $n$ $R^{2}$ $\alpha$ $\log_{10}\left(STR_{(c,t)}\right)$ $\log_{10}\left(POP_{(c,t)}\right)$ $\log_{10}\left(SGDP_{(c,t)}\right)$ $\log_{10}\left(POL_{(c,t)}\right)$ $t=1$ $t=2$ $t=3$ $t=4$ $t=5$ $t=6$ $t=7$ $t=8$ Dependent variable: $\log_{10}\left(CON_{(c,t+1)}\right)$ P1TD 141 0.38 -1.40*** (0.25) 0.43*** (0.12) -0.03 (0.13) 0.03 (0.16) 0.07 (0.15) 0.34* (0.17) 0.34 (0.18) 0.41* (0.19) 0.46* (0.19) 0.54** (0.17) P2TD 141 0.48 -4.16*** (1.24) 0.44*** (0.11) 0.40* (0.19) -0.06 (0.15) -0.03 (0.16) 0.06 (0.16) 0.28 (0.16) 0.28 (0.17) 0.34 (0.18) 0.39* (0.18) 0.48** (0.17) P3TD 141 0.55 -7.88*** (2.00) 0.34* (0.14) 0.39* (0.17) 1.08* (0.42) -0.09 (0.14) -0.04 (0.16) 0.05 (0.16) 0.25 (0.15) 0.24 (0.16) 0.31 (0.17) 0.37* (0.16) 0.46** (0.16) P4TD 132 0.56 -11.06*** (2.41) 0.42** (0.15) 0.38* (0.16) 1.26*** (0.38) 0.97 (0.61) -0.11 (0.14) -0.04 (0.16) 0.02 (0.16) 0.20 (0.16) 0.19 (0.17) 0.31 (0.19) 0.36 (0.19) 0.41* (0.19) Dependent variable: $\log_{10}\left(CON_{(c,t+2)}\right)$ P1TD 147 0.38 -1.45*** (0.26) 0.44*** (0.13) 0.09 (0.11) 0.09 (0.12) 0.40* (0.17) 0.39* (0.18) 0.46* (0.18) 0.52** (0.19) 0.58** (0.18) 0.49** (0.18) P2TD 147 0.49 -4.50*** (1.20) 0.46*** (0.12) 0.44* (0.19) 0.04 (0.10) 0.11 (0.12) 0.36* (0.15) 0.35* (0.16) 0.41** (0.16) 0.47** (0.17) 0.54*** (0.16) 0.45** (0.16) P3TD 147 0.59 -8.48*** (1.93) 0.35* (0.14) 0.41* (0.17) 1.17** (0.42) 0.06 (0.11) 0.11 (0.13) 0.33* (0.15) 0.32* (0.16) 0.39* (0.16) 0.46** (0.17) 0.54** (0.16) 0.45** (0.16) P4TD 137 0.60 -11.92*** (2.21) 0.43** (0.14) 0.42** (0.16) 1.36*** (0.36) 1.05 (0.60) 0.07 (0.12) 0.08 (0.13) 0.28 (0.16) 0.26 (0.17) 0.34 (0.18) 0.46* (0.20) 0.48* (0.20) 0.44* (0.19) Dependent variable: $\log_{10}\left(CON_{(c,t+3)}\right)$ P1TD 135 0.34 -1.33*** (0.24) 0.43*** (0.13) 0.05 (0.10) 0.31* (0.15) 0.32* (0.15) 0.40** (0.15) 0.46** (0.16) 0.53*** (0.15) 0.42** (0.16) P2TD 135 0.45 -4.32*** (1.28) 0.44*** (0.12) 0.42* (0.20) 0.10 (0.11) 0.31* (0.13) 0.33* (0.13) 0.39** (0.13) 0.46** (0.14) 0.53*** (0.13) 0.41** (0.13) P3TD 135 0.57 -8.48*** (1.97) 0.34* (0.14) 0.39* (0.18) 1.24** (0.45) 0.06 (0.11) 0.25* (0.12) 0.25* (0.12) 0.32** (0.11) 0.39** (0.13) 0.48*** (0.12) 0.39** (0.13) P4TD 128 0.62 -12.69*** (2.08) 0.46*** (0.14) 0.38* (0.16) 1.48*** (0.34) 1.34* (0.55) 0.02 (0.10) 0.15 (0.12) 0.15 (0.12) 0.23 (0.12) 0.30* (0.13) 0.39** (0.15) 0.32* (0.15) Dependent variable: $\log_{10}\left(CON_{(c,t+4)}\right)$ P1TD 125 0.28 -1.22*** (0.25) 0.39** (0.13) 0.28 (0.16) 0.25 (0.18) 0.37* (0.18) 0.43* (0.19) 0.50** (0.19) 0.40* (0.19) P2TD 125 0.38 -4.02** (1.40) 0.41*** (0.12) 0.40 (0.21) 0.23 (0.14) 0.20 (0.16) 0.31 (0.16) 0.37* (0.17) 0.45** (0.17) 0.34* (0.17) P3TD 125 0.53 -8.57*** (2.07) 0.30* (0.14) 0.37 (0.20) 1.35** (0.47) 0.19 (0.13) 0.14 (0.15) 0.24 (0.15) 0.32* (0.16) 0.41* (0.17) 0.33* (0.16) P4TD 121 0.57 -11.87*** (1.98) 0.41** (0.14) 0.34* (0.17) 1.46*** (0.36) 1.18* (0.57) 0.15 (0.14) 0.08 (0.18) 0.19 (0.18) 0.26 (0.18) 0.34 (0.20) 0.30 (0.20) Dependent variable: $\log_{10}\left(CON_{(c,t+5)}\right)$ P1TD 113 0.22 -0.86** (0.29) 0.36** (0.13) -0.02 (0.10) 0.07 (0.09) 0.16 (0.10) 0.22* (0.10) 0.13 (0.09) P2TD 113 0.34 -3.76** (1.43) 0.37** (0.12) 0.41 (0.22) -0.02 (0.09) 0.06 (0.09) 0.15 (0.10) 0.22* (0.10) 0.12 (0.09) P3TD 113 0.52 -8.59*** (2.09) 0.26 (0.14) 0.37 (0.20) 1.43** (0.50) -0.05 (0.09) 0.02 (0.08) 0.12 (0.10) 0.20* (0.10) 0.13 (0.08) P4TD 111 0.58 -12.01*** (2.02) 0.39** (0.14) 0.32* (0.16) 1.53*** (0.37) 1.28* (0.61) -0.05 (0.10) 0.01 (0.10) 0.10 (0.10) 0.17 (0.11) 0.13 (0.09) Dependent variable: $\log_{10}(CON_{\left(c,t+1:3\right)})$ P1TD 122 0.40 -1.47*** (0.25) 0.44*** (0.12) -0.03 (0.07) 0.15 (0.10) 0.36** (0.12) 0.40* (0.20) 0.44* (0.20) 0.54** (0.18) 0.56** (0.19) P2TD 122 0.53 -4.85*** (1.27) 0.45*** (0.12) 0.48* (0.19) -0.03 (0.07) 0.13 (0.09) 0.37** (0.13) 0.36* (0.17) 0.40* (0.17) 0.50** (0.15) 0.52*** (0.15) P3TD 122 0.59 -8.30*** (2.04) 0.36** (0.13) 0.48** (0.18) 0.98* (0.39) -0.04 (0.07) 0.13 (0.10) 0.35* (0.15) 0.33 (0.17) 0.38* (0.17) 0.49** (0.16) 0.51** (0.16) P4TD 116 0.58 -9.54*** (2.74) 0.39* (0.16) 0.48** (0.18) 1.05* (0.41) 0.38 (0.68) -0.05 (0.07) 0.12 (0.11) 0.34* (0.15) 0.31 (0.18) 0.36* (0.18) 0.49** (0.19) 0.50** (0.19) Dependent variable: $\log_{10}(CON_{\left(c,t+1:5\right)})$ P1TD 80 0.47 -1.30*** (0.20) 0.43*** (0.08) 0.11* (0.05) 0.18* (0.07) 0.32*** (0.08) 0.38* (0.15) 0.41** (0.15) P2TD 80 0.58 -3.89*** (1.11) 0.43*** (0.08) 0.37* (0.17) 0.10* (0.05) 0.17* (0.07) 0.34*** (0.09) 0.36** (0.14) 0.39** (0.13) P3TD 80 0.69 -7.95*** (1.72) 0.34*** (0.10) 0.38** (0.14) 1.13*** (0.31) 0.05 (0.06) 0.11 (0.08) 0.30** (0.10) 0.31* (0.13) 0.35** (0.12) P4TD 80 0.71 -10.38*** (2.07) 0.42*** (0.12) 0.38** (0.13) 1.26*** (0.26) 0.74 (0.50) 0.04 (0.06) 0.10 (0.08) 0.28** (0.10) 0.26 (0.13) 0.30* (0.13)
@@
Table 5: Pooled regression results with year-specific dummy variables. We measure $t$ relative to 2006 (the start of our data collection period). Thus, $\textit{t}=1$” denotes a dummy variable for 2007, $\textit{t}=2$” denotes a dummy variable for 2008, and so on. The table contains results for multiple model specifications, numbered P1TD to P4TD, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level; and *** at the $0.001$ $p$-level.
@@ 5.2 Fixed Effects Regression Results
Table 6 presents our fixed effects regression results. As before, the table contains results for multiple model specifications, now numbered F1 to F4, and multiple dependent variables as explained in Section 4. For illustration purposes, Figure 4 depicts our most basic fixed effects regression model (F1) without any control variables.
@@
F1 113 0.02 24.66*** 0.19 (0.26)
F2 113 0.06 21.14*** 0.18 (0.26) 8.39 (7.06)
F3 113 0.07 14.49*** 0.19 (0.26) 9.17 (7.31) 1.39 (1.41)
F4 111 0.09 12.20*** 0.24 (0.29) 8.40 (6.88) 1.51 (1.60) -1.01 (0.81)
Dependent variable: $\log_{10}(CON_{\left(c,t+1:3\right)})$
F1 122 0.22 22.22*** 0.76** (0.26)
F2 122 0.22 16.58*** 0.77** (0.26) 3.93 (11.31)
F3 122 0.22 14.04*** 0.78** (0.25) 4.19 (11.18) 0.32 (1.05)
F4 116 0.26 14.49*** 0.76** (0.25) 5.40 (11.78) 0.56 (1.10) -1.78 (1.41)
Dependent variable: $\log_{10}(CON_{\left(c,t+1:5\right)})$
F1 80 0.14 17.23*** 0.53* (0.25)
F2 80 0.18 14.09*** 0.53* (0.26) 8.68 (13.63)
F3 80 0.23 10.68*** 0.60* (0.23) 9.38 (12.89) 3.05** (0.98)
F4 80 0.30 10.15*** 0.58** (0.21) 10.42 (12.95) 3.05*** (0.77) -2.14 (1.72)
Table 6: Fixed effects regression results. The table contains results for multiple model specifications, numbered F1 to F4, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level.
All models in Table 6 have significant F-statistics for poolability, indicating that a fixed effects regression model is appropriate (and that a pooled regression model is not). Our $\log_{10}(STR_{(c,t)})$ coefficients are statistically significant except when we use $\log_{10}(CON_{(c,t+4)})$ or $\log_{10}(CON_{(c,t+5)})$ as the dependent variable, indicating that the relationship between STRs and convictions might disappear after three years. When significant, the fixed effects $\log_{10}(STR_{(c,t)})$ coefficients are generally substantially larger than the pooled $\log_{10}(STR_{(c,t)})$ coefficients (compare Tables 3 and 6), indicating that the relationship between STRs and convictions is affected by cross-country differences. Our largest $\log_{10}(STR_{(c,t)})$ coefficient equals $0.80$ with a standard error of $0.19$ and a $95\%$ confidence interval of $[0.42:1.18]$. While the confidence interval allows a coefficient larger than 1, our results consistently suggest a sub-linear relationship between STRs and convictions.
Fig. 4: Illustration of our most basic fixed effects regression model (F1). $STR_{(c,t)}$ denotes the number of STRs filed in country $c$ in year $t$ per 100,000 capita. $CON_{(c,t+1:3)}$ denotes the average number of money laundering convictions per year in country $c$ up to three years after $t$ per 100,000 capita. We measure $t$ relative to 2006 (the start of our data collection). Thus, $\textit{t}=0$ corresponds to 2006, $\textit{t}=1$ corresponds to 2007, and so on.
@@
Table 7: Estimates of country fixed effects $\alpha_{(c)}$ and $A_{(c)}$ using model F1 and dependent variable $\log_{10}(CON_{(c,t+1:3)})$. The table should not be seen as a ranking; there are many reasons why countries might have different $\alpha_{(c)}$ and $A_{(c)}$ values. Values are impacted by our scaling.
To complement Figure 4, Table 7 presents estimates of the country-specific fixed effects $A_{(c)}$ and $\alpha_{(c)}$ estimated in model F1 with $\log_{10}(CON_{(c,t+1:3)})$ as the dependent variable (i.e., the model depicted in Figure 4). Recall that the effects measure a country’s “scale of production;” see equations (4) and (5). Note that $\alpha_{(c)}$ is measured on a $\log_{10}$-scale (why it can be negative). The effects may reflect different legal frameworks, crime types, prosecutorial priorities, or how financial intelligence units work. In particular, the variety in the estimated fixed effects could indicate that AML regulation and enforcement is not homogenized across EU countries. We stress that the table should not be seen as a ranking. Spain clearly stands out in the table; likely an indication that it handles (or counts) money laundering cases differently. [^19] This observation makes the use of a fixed effects model more appropriate.
@@
F1T 113 0.27 33.54*** -0.09 (0.20) 0.07*** (0.02)
F2T 113 0.27 27.73*** -0.08 (0.21) 1.91 (4.01) 0.07*** (0.02)
F3T 113 0.27 18.88*** -0.08 (0.21) 1.95 (4.51) 0.05 (1.84) 0.07*** (0.02)
F4T 111 0.33 17.17*** -0.03 (0.21) 0.15 (3.66) 0.04 (2.01) -1.62** (0.59) 0.07*** (0.02)
Dependent variable: $\log_{10}(CON_{\left(c,t+1:3\right)})$
F1T 122 0.52 33.03*** 0.18 (0.16) 0.09*** (0.02)
F2T 122 0.52 24.65*** 0.17 (0.17) -1.23 (5.49) 0.09*** (0.02)
F3T 122 0.57 23.52*** -0.04 (0.18) -5.52 (5.66) -4.26** (1.27) 0.10*** (0.02)
F4T 116 0.59 24.37*** -0.01 (0.18) -3.91 (6.22) -4.09** (1.38) -1.37 (0.92) 0.10*** (0.02)
Dependent variable: $\log_{10}(CON_{\left(c,t+1:5\right)})$
F1T 80 0.63 37.39*** 0.16 (0.11) 0.09*** (0.02)
F2T 80 0.63 29.42*** 0.17 (0.10) 3.46 (6.06) 0.09*** (0.02)
F3T 80 0.64 21.50*** 0.10 (0.11) 2.65 (6.25) -1.63 (1.19) 0.09*** (0.02)
F4T 80 0.71 25.04*** 0.08 (0.08) 3.68 (5.87) -1.66 (1.24) -2.24 (1.16) 0.09*** (0.01)
Table 8: Fixed effects regression results with $t$ (i.e., “time”) included as a regressor. The table contains results for multiple model specifications, F1T to F4T, and multiple dependent variables. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level.

### v2

@@ 5.1 Pooled Regression Results
Table 4: Pooled regression results with $t$ (i.e., “time”) as a regressor. The table contains results for multiple model specifications, named PT1 to PT4, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level; and *** at the $0.001$ $p$-level.
Model ID $n$ $R^{2}$ $\alpha$ $\log_{10}\left(STR_{(c,t)}\right)$ $\log_{10}\left(POP_{(c,t)}\right)$ $\log_{10}\left(SGDP_{(c,t)}\right)$ $\log_{10}\left(POL_{(c,t)}\right)$ $t=1$ $t=2$ $t=3$ $t=4$ $t=5$ $t=6$ $t=7$ $t=8$ Dependent variable: $\log_{10}\left(CON_{(c,t+1)}\right)$ PTD1 141 0.38 -1.40*** (0.25) 0.43*** (0.12) -0.03 (0.13) 0.03 (0.16) 0.07 (0.15) 0.34* (0.17) 0.34 (0.18) 0.41* (0.19) 0.46* (0.19) 0.54** (0.17) PTD2 141 0.48 -4.16*** (1.24) 0.44*** (0.11) 0.40* (0.19) -0.06 (0.15) -0.03 (0.16) 0.06 (0.16) 0.28 (0.16) 0.28 (0.17) 0.34 (0.18) 0.39* (0.18) 0.48** (0.17) PTD3 141 0.55 -7.88*** (2.00) 0.34* (0.14) 0.39* (0.17) 1.08* (0.42) -0.09 (0.14) -0.04 (0.16) 0.05 (0.16) 0.25 (0.15) 0.24 (0.16) 0.31 (0.17) 0.37* (0.16) 0.46** (0.16) PTD4 132 0.56 -11.06*** (2.41) 0.42** (0.15) 0.38* (0.16) 1.26*** (0.38) 0.97 (0.61) -0.11 (0.14) -0.04 (0.16) 0.02 (0.16) 0.20 (0.16) 0.19 (0.17) 0.31 (0.19) 0.36 (0.19) 0.41* (0.19) Dependent variable: $\log_{10}\left(CON_{(c,t+2)}\right)$ PTD1 147 0.38 -1.45*** (0.26) 0.44*** (0.13) 0.09 (0.11) 0.09 (0.12) 0.40* (0.17) 0.39* (0.18) 0.46* (0.18) 0.52** (0.19) 0.58** (0.18) 0.49** (0.18) PTD2 147 0.49 -4.50*** (1.20) 0.46*** (0.12) 0.44* (0.19) 0.04 (0.10) 0.11 (0.12) 0.36* (0.15) 0.35* (0.16) 0.41** (0.16) 0.47** (0.17) 0.54*** (0.16) 0.45** (0.16) PTD3 147 0.59 -8.48*** (1.93) 0.35* (0.14) 0.41* (0.17) 1.17** (0.42) 0.06 (0.11) 0.11 (0.13) 0.33* (0.15) 0.32* (0.16) 0.39* (0.16) 0.46** (0.17) 0.54** (0.16) 0.45** (0.16) PTD4 137 0.60 -11.92*** (2.21) 0.43** (0.14) 0.42** (0.16) 1.36*** (0.36) 1.05 (0.60) 0.07 (0.12) 0.08 (0.13) 0.28 (0.16) 0.26 (0.17) 0.34 (0.18) 0.46* (0.20) 0.48* (0.20) 0.44* (0.19) Dependent variable: $\log_{10}\left(CON_{(c,t+3)}\right)$ PTD1 135 0.34 -1.33*** (0.24) 0.43*** (0.13) 0.05 (0.10) 0.31* (0.15) 0.32* (0.15) 0.40** (0.15) 0.46** (0.16) 0.53*** (0.15) 0.42** (0.16) PTD2 135 0.45 -4.32*** (1.28) 0.44*** (0.12) 0.42* (0.20) 0.10 (0.11) 0.31* (0.13) 0.33* (0.13) 0.39** (0.13) 0.46** (0.14) 0.53*** (0.13) 0.41** (0.13) PTD3 135 0.57 -8.48*** (1.97) 0.34* (0.14) 0.39* (0.18) 1.24** (0.45) 0.06 (0.11) 0.25* (0.12) 0.25* (0.12) 0.32** (0.11) 0.39** (0.13) 0.48*** (0.12) 0.39** (0.13) PTD4 128 0.62 -12.69*** (2.08) 0.46*** (0.14) 0.38* (0.16) 1.48*** (0.34) 1.34* (0.55) 0.02 (0.10) 0.15 (0.12) 0.15 (0.12) 0.23 (0.12) 0.30* (0.13) 0.39** (0.15) 0.32* (0.15) Dependent variable: $\log_{10}\left(CON_{(c,t+4)}\right)$ PTD1 125 0.28 -1.22*** (0.25) 0.39** (0.13) 0.28 (0.16) 0.25 (0.18) 0.37* (0.18) 0.43* (0.19) 0.50** (0.19) 0.40* (0.19) PTD2 125 0.38 -4.02** (1.40) 0.41*** (0.12) 0.40 (0.21) 0.23 (0.14) 0.20 (0.16) 0.31 (0.16) 0.37* (0.17) 0.45** (0.17) 0.34* (0.17) PTD3 125 0.53 -8.57*** (2.07) 0.30* (0.14) 0.37 (0.20) 1.35** (0.47) 0.19 (0.13) 0.14 (0.15) 0.24 (0.15) 0.32* (0.16) 0.41* (0.17) 0.33* (0.16) PTD4 121 0.57 -11.87*** (1.98) 0.41** (0.14) 0.34* (0.17) 1.46*** (0.36) 1.18* (0.57) 0.15 (0.14) 0.08 (0.18) 0.19 (0.18) 0.26 (0.18) 0.34 (0.20) 0.30 (0.20) Dependent variable: $\log_{10}\left(CON_{(c,t+5)}\right)$ PTD1 113 0.22 -0.86** (0.29) 0.36** (0.13) -0.02 (0.10) 0.07 (0.09) 0.16 (0.10) 0.22* (0.10) 0.13 (0.09) PTD2 113 0.34 -3.76** (1.43) 0.37** (0.12) 0.41 (0.22) -0.02 (0.09) 0.06 (0.09) 0.15 (0.10) 0.22* (0.10) 0.12 (0.09) PTD3 113 0.52 -8.59*** (2.09) 0.26 (0.14) 0.37 (0.20) 1.43** (0.50) -0.05 (0.09) 0.02 (0.08) 0.12 (0.10) 0.20* (0.10) 0.13 (0.08) PTD4 111 0.58 -12.01*** (2.02) 0.39** (0.14) 0.32* (0.16) 1.53*** (0.37) 1.28* (0.61) -0.05 (0.10) 0.01 (0.10) 0.10 (0.10) 0.17 (0.11) 0.13 (0.09) Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$ PTD1 122 0.40 -1.47*** (0.25) 0.44*** (0.12) -0.03 (0.07) 0.15 (0.10) 0.36** (0.12) 0.40* (0.20) 0.44* (0.20) 0.54** (0.18) 0.56** (0.19) PTD2 122 0.53 -4.85*** (1.27) 0.45*** (0.12) 0.48* (0.19) -0.03 (0.07) 0.13 (0.09) 0.37** (0.13) 0.36* (0.17) 0.40* (0.17) 0.50** (0.15) 0.52*** (0.15) PTD3 122 0.59 -8.30*** (2.04) 0.36** (0.13) 0.48** (0.18) 0.98* (0.39) -0.04 (0.07) 0.13 (0.10) 0.35* (0.15) 0.33 (0.17) 0.38* (0.17) 0.49** (0.16) 0.51** (0.16) PTD4 116 0.58 -9.54*** (2.74) 0.39* (0.16) 0.48** (0.18) 1.05* (0.41) 0.38 (0.68) -0.05 (0.07) 0.12 (0.11) 0.34* (0.15) 0.31 (0.18) 0.36* (0.18) 0.49** (0.19) 0.50** (0.19) Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$ PTD1 80 0.47 -1.30*** (0.20) 0.43*** (0.08) 0.11* (0.05) 0.18* (0.07) 0.32*** (0.08) 0.38* (0.15) 0.41** (0.15) PTD2 80 0.58 -3.89*** (1.11) 0.43*** (0.08) 0.37* (0.17) 0.10* (0.05) 0.17* (0.07) 0.34*** (0.09) 0.36** (0.14) 0.39** (0.13) PTD3 80 0.69 -7.95*** (1.72) 0.34*** (0.10) 0.38** (0.14) 1.13*** (0.31) 0.05 (0.06) 0.11 (0.08) 0.30** (0.10) 0.31* (0.13) 0.35** (0.12) PTD4 80 0.71 -10.38*** (2.07) 0.42*** (0.12) 0.38** (0.13) 1.26*** (0.26) 0.74 (0.50) 0.04 (0.06) 0.10 (0.08) 0.28** (0.10) 0.26 (0.13) 0.30* (0.13)
@@
Table 5: Pooled regression results with year-specific dummy variables. We measure $t$ relative to 2006 (the start of our data collection period). Thus, $\textit{t}=1$” denotes a dummy variable for 2007, $\textit{t}=2$” denotes a dummy variable for 2008, and so on. The table contains results for multiple model specifications, named PTD1 to PTD4, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level; and *** at the $0.001$ $p$-level.
@@ 5.2 Fixed Effects Regression Results
Table 6 presents our fixed effects regression results. As before, the table contains results for multiple model specifications, now named FE1 to FE4, and multiple dependent variables as explained in Section 4. For illustration purposes, Figure 4 depicts our most basic fixed effects regression model (F1) without any control variables. A Q-Q plot of residuals is displayed in Appendix A.
@@
FE1 113 0.02 24.66*** 0.19 (0.26)
FE2 113 0.06 21.14*** 0.18 (0.26) 8.39 (7.06)
FE3 113 0.07 14.49*** 0.19 (0.26) 9.17 (7.31) 1.39 (1.41)
FE4 111 0.09 12.20*** 0.24 (0.29) 8.40 (6.88) 1.51 (1.60) -1.01 (0.81)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$
FE1 122 0.22 22.22*** 0.76** (0.26)
FE2 122 0.22 16.58*** 0.77** (0.26) 3.93 (11.31)
FE3 122 0.22 14.04*** 0.78** (0.25) 4.19 (11.18) 0.32 (1.05)
FE4 116 0.26 14.49*** 0.76** (0.25) 5.40 (11.78) 0.56 (1.10) -1.78 (1.41)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$
FE1 80 0.14 17.23*** 0.53* (0.25)
FE2 80 0.18 14.09*** 0.53* (0.26) 8.68 (13.63)
FE3 80 0.23 10.68*** 0.60* (0.23) 9.38 (12.89) 3.05** (0.98)
FE4 80 0.30 10.15*** 0.58** (0.21) 10.42 (12.95) 3.05*** (0.77) -2.14 (1.72)
Table 6: Fixed effects regression results. The table contains results for multiple model specifications, named FE1 to FE4, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level.
All models in Table 6 have significant F-statistics for poolability, indicating that fixed effects regression models are appropriate (and that pooled regression models are not). Our $\log_{10}(STR_{(c,t)})$ coefficients are statistically significant except when we use $\log_{10}(CON_{(c,t+4)})$ or $\log_{10}(CON_{(c,t+5)})$ as the dependent variable, indicating that the relationship between STRs and convictions might disappear after three years. When significant, the fixed effects $\log_{10}(STR_{(c,t)})$ coefficients are generally substantially larger than the pooled $\log_{10}(STR_{(c,t)})$ coefficients (compare Tables 3 and 6), indicating that the relationship between STRs and convictions is affected by cross-country differences. Our largest $\log_{10}(STR_{(c,t)})$ coefficient equals $0.80$ with a standard error of $0.19$ and a $95\%$ confidence interval of $[0.42:1.18]$. While the confidence interval allows a coefficient larger than 1, our results consistently suggest a sub-linear relationship between STRs and convictions.
Fig. 4: Illustration of our most basic fixed effects regression model (FE1). $STR_{(c,t)}$ denotes the number of STRs filed in country $c$ in year $t$ per 100,000 capita. $CON_{(c,t+1:3)}$ denotes the average number of money laundering convictions per year in country $c$ up to three years after $t$ per 100,000 capita. We measure $t$ relative to 2006 (the start of our data collection). Thus, $\textit{t}=0$ corresponds to 2006, $\textit{t}=1$ corresponds to 2007, and so on.
@@
Table 7: Estimates of country fixed effects $\alpha_{(c)}$ and $A_{(c)}$ using model FE1 and dependent variable $\log_{10}(CON_{(c,t+1:3)})$. The table should not be seen as a ranking; there are many reasons why countries might have different $\alpha_{(c)}$ and $A_{(c)}$ values. Values are impacted by our scaling.
To complement Figure 4, Table 7 presents estimates of the country-specific fixed effects $A_{(c)}$ and $\alpha_{(c)}$ estimated in model FE1 with $\log_{10}(CON_{(c,t+1:3)})$ as the dependent variable (i.e., the model depicted in Figure 4). Recall that the effects measure a country’s “scale of production;” see equations (4) and (5). Note that $\alpha_{(c)}$ is measured on a $\log_{10}$-scale (why it can be negative). The effects may reflect different legal frameworks, crime types, prosecutorial priorities, or how financial intelligence units work. In particular, the variety in the estimated fixed effects could indicate that AML regulation and enforcement is not homogenized across EU countries. We stress that the table should not be seen as a ranking. Spain clearly stands out in the table; likely an indication that it handles (or counts) money laundering cases differently. [^20] This observation makes the use of a fixed effects model more appropriate.
@@
FET1 113 0.27 33.54*** -0.09 (0.20) 0.07*** (0.02)
FET2 113 0.27 27.73*** -0.08 (0.21) 1.91 (4.01) 0.07*** (0.02)
FET3 113 0.27 18.88*** -0.08 (0.21) 1.95 (4.51) 0.05 (1.84) 0.07*** (0.02)
FET4 111 0.33 17.17*** -0.03 (0.21) 0.15 (3.66) 0.04 (2.01) -1.62** (0.59) 0.07*** (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$
FET1 122 0.52 33.03*** 0.18 (0.16) 0.09*** (0.02)
FET2 122 0.52 24.65*** 0.17 (0.17) -1.23 (5.49) 0.09*** (0.02)
FET3 122 0.57 23.52*** -0.04 (0.18) -5.52 (5.66) -4.26** (1.27) 0.10*** (0.02)
FET4 116 0.59 24.37*** -0.01 (0.18) -3.91 (6.22) -4.09** (1.38) -1.37 (0.92) 0.10*** (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$
FET1 80 0.63 37.39*** 0.16 (0.11) 0.09*** (0.02)
FET2 80 0.63 29.42*** 0.17 (0.10) 3.46 (6.06) 0.09*** (0.02)
FET3 80 0.64 21.50*** 0.10 (0.11) 2.65 (6.25) -1.63 (1.19) 0.09*** (0.02)
FET4 80 0.71 25.04*** 0.08 (0.08) 3.68 (5.87) -1.66 (1.24) -2.24 (1.16) 0.09*** (0.01)
Table 8: Fixed effects regression results with $t$ (i.e., “time”) included as a regressor. The table contains results for multiple model specifications, named FET1 to FET4, and multiple dependent variables. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level.